# Beyond Community Notes: A Framework for Understanding and Building Crowdsourced Context Systems for Social Media

### v2

@@ 5.1. Systematic Literature Review
Figure 3. PRISMA flow diagram (pagePRISMA2020Statement2021) for our systematic literature review, showing both the database search and the backwards reference search. We analyzed a final set of papers that combined the results of these two searches (n=56).:
In order to synthesize the research community’s understanding of Crowdsourced Context Systems, we performed a PRISMA-style systematic literature review (pagePRISMA2020Statement2021). We sourced relevant papers using both a database keyword search and a backwards reference search (jalaliSystematicLiteratureStudies2012). Figure 3 shows a flow diagram summarizing our search and screening process. Our goal was to identify original research that directly studies CCS implementations. Since this category of system has not been consistently named prior to our work, we focused our review on studies of the first large-scale implementation of a CCS by a major social media platform: Twitter’s Community Notes/Birdwatch. We thus used “Birdwatch” and “Community Notes” as keywords and restricted our search to articles that have been written since 2021, when Twitter first launched their implementation. Note that since this is an emerging technology and the set of articles is limited, we decided to include sources of knowledge that are valued in the HCI community but are sometimes excluded from literature reviews, such as pre-prints and non-full-length conference submissions. We thus settled on the following inclusion and exclusion criteria:
@@
Exclusion criteria: The article was not written in English, does not present original research, is only an abstract/a summary of other research, or is not about the functioning of a Twitter-style CCS.
@@
We conducted our database search in November 2025. We searched seven databases that are well-known in the fields of HCI and social media research. We began with six that were used in another recent HCI literature review of social media research (zhangWhatWeMean2023): ACM Digital Library, IEEE Xplore, SAGE Journals, ScienceDirect (Elsevier), Taylor & Francis Online (Tandfonline), Emerald Insight. We additionally searched arXiv in order to also include preprints, which are common in these fields.
We searched the full-text, title, abstract, and metadata of articles using the following case-insensitive query string, and limit our results to articles published in 2021 or later:
@@
This search resulted in n=590 results.
@@
We screened the search results by first considering their title. One researcher read the title for each result and removed articles that were either duplicates or were obviously off-topic, such as those that were not about social media systems. This resulted in n=254 articles.
@@
One researcher read the abstracts of the remaining articles to further screen the results. We removed redundant articles that only summarized research that was contained elsewhere in our sample, such as books and book chapters, editorials, abstracts, keynotes, and complete conference proceedings. We also removed articles that were topically irrelevant, such as those that did not focus on a Twitter-style CCS, or that used data from such a system, but only to answer research questions unrelated to the system. This resulted in n=47 articles.
@@
Two researchers then read the full-text of the remaining 47 articles and applied the inclusion and exclusion criteria to screen more thoroughly. The researchers met during the process to compare their results and discuss any differences or challenging cases until they were in agreement. This produced a set of n=40 articles.
@@
The first author then used this set of 40 articles as the starting set for a backwards reference search and extracted all of the n=2,207 references included in these articles. We de-duplicated these references by paper title, which resulted in a set of n=1,385 unique papers. We then performed the same screening steps that we used in the forward search: one researcher first screened the titles of the cited papers, which resulted in n=175 papers, then screened this set by abstract, which resulted in n=23, then finally two researchers performed a full-text screen. At the end of this screening process we had identified n=16 relevant papers that were not found in our forward search. We combined the sets from the two searches and produced a final set of n=56 papers.
@@
Two researchers then agreed upon data to extract from each article in the final set, which included the data source, the research questions, the main findings, and any design implications mentioned. The two researchers independently read the final set of articles and extracted this information. They then compared values, discussed any areas of disagreement, and came to a consensus for the extracted values for each paper.
Next the two researchers independently grouped similar papers into affinity groups (beyer99) based on the primary focus of the study. The researchers reconvened, compared their groups, and discussed the reasons for their groupings. Several papers were shifted between groups based on the discussion. Once the researchers agreed upon the groups, they came up with labels to describe each group based on the focus of the studies (see Table 1).
@@ 5.2. Design Feature Identification
In order to identify the design decisions that shape CCS, two authors analyzed four real-world implementations of CCS: Twitter/X’s Community Notes [^3] https://x.com/i/communitynotes/, Meta’s Community Notes [^4] https://transparency.meta.com/features/community-notes (which covers Facebook, Instagram, and Threads), TikTok’s Footnotes [^5] https://support.TikTok.com/en/using-TikTok/exploring-videos/footnotes, and YouTube’s Notes [^6] https://support.google.com/YouTube/answer/14925346?hl=en. Our analysis is based on public, company-written documentation and communications about the systems and their features, which we will collectively refer to as articles. In addition, we had access to some of the systems, which we used to complement our understanding. Unfortunately, while we attempted to become Contributors for all of the platforms, some (TikTok, YouTube) did not grant us access.
The researchers independently read through the articles one at a time and extracted design details into a spreadsheet, which resulted in a set of design features for each platform. After independently reviewing the articles for a platform, the researchers convened and performed a round of affinity diagramming (beyer99), in which they wrote the design features on sticky notes and placed them on a whiteboard. They placed similar design features together, and through this inductive method identified feature groups. The researchers then performed thematic analysis (braunUsingThematicAnalysis2006) to identify the themes of the groupings and came up with provisional group names. They repeated this process with the articles for each platform, reconvening to do another round of affinity diagramming after each round of feature extraction. These design features and feature groups were further synthesized into the aspects, dimensions, and options of the design space, which we present below in Section 7.
@@ 6. What Does the Research Community Know About Twitter-Style CCS?
Study Focus Primary Method Papers
System Impact (n=24) Data Analysis: X Community Notes Data (bouchaudAlgorithmicResolutionCrowdsourced2025b; renaultCollaborativelyAddingContext2024a; chuai25; truong2025communitynotesvulnerablerater; slaughter25; chujyo25; chuaiCommunitybasedFactcheckingReduces2024a; rollout21; kimDifferentialImpactIndividual2025e; diffusion21; chuai2025requestnoterequestfunction; matamoros25; razuvayevskaya2025timelinessconsensuscompositioncrowd; kangur2024checkscheckersexploringsource; DeepDiveXs; gao24; jama24; bobek2025communityfactchecksbreakfollower; borwankarDemocratizationMisinformationMonitoring2022a; Wirtschafter_Majumder_2023; wangEfficiencyCommunityBasedContent2024b; colemanLimitingFactorsEffectiveness2023; dekeulenaarTwitterDemotionCommunity2025; cham)
Contributor Behavior (n=16) Data Analysis: X Community Notes Data (allen22; borenstein2025; pilarskiCommunityNotesVs2023; pröllochs2021communitybasedfactcheckingtwittersbirdwatch; cikm22; kuuse25; jones22; solovevReferencesUnbiasedSources2025a; arjmandilari2025threatssustainabilitycommunitynotes; toyoda25; Phillips2025; renault25; fraxanetUnpackingPolarizationAntagonism2024c; martelPoliticalMotivesHelp2025; mohammadi2025birdwatchcommunitynotestwitter; yoon25)
System Design (n=10) User studies, design specs, algorithm analysis (mohammadiAIFeedbackEnhances2025a; wu2025crowdllmaugmentedcommunitynotes; wojcikBirdwatchCrowdWisdom2022b; hawkeye; franzmeyer2024hellofreshllmevaluationsstreams; liScalingHumanJudgment2025a; supernotes; zhang2025commenotessynthesizingorganiccomments; singh2025on; xing2025communitynotesdatasetexploringhelpfulness)
Users’ Perceptions (n=6) User studies (sharevski2025helps; liu23; shusas24; drolsbachCommunityNotesIncrease2024a; Kankham18062025; Hameleers18052024)
Table 1. The n=56 articles found in our literature review, grouped by study focus.:
@@
Our systematic literature review of research into Twitter-style CCS found n=56 papers and identified four high-level categories of study: studies of the system’s impact, contributor behavior, alternative system designs, and users’ perceptions of crowdsourced context. These categories naturally overlap and are not mutually exclusive, but we highlight them in order to emphasize the different types of knowledge emerging from this body of research. While some papers make multiple contributions and thus span categories, we present them in Table 1 grouped according to their most significant contribution for simplicity.
@@
The most common type of study explored the impact of a CCS on its implementing platform (n=24 papers). These studies all performed analysis of data from X’s Community Notes. These studies asked both causal questions, such as about the effect that displaying a note has on post diffusion (diffusion21; rollout21), as well as descriptive questions about the performance of X’s system in several specific use cases (chujyo25; bouchaudAlgorithmicResolutionCrowdsourced2025b). The studies found strong evidence that CCS can both reduce the spread (diffusion21) and believability (wojcikBirdwatchCrowdWisdom2022b) of misleading posts. The research also argued that the bridging-based algorithm in X’s Community Notes system has demonstrated it can identify high-quality crowdsourced notes that are likely to be perceived as helpful by users across an ideological spectrum (bouchaudAlgorithmicResolutionCrowdsourced2025b). However, the studies also showed that this bridging approach takes a long time to identify “helpful” notes for misleading posts. As a result, there is often a significant delay before notes are displayed to users (renaultCollaborativelyAddingContext2024a; chuaiCommunitybasedFactcheckingReduces2024a; rollout21). Additionally, studies have shown that the vast majority of suggested notes ($\sim 90\%$) are never published due to either a lack of ratings or consensus  (chujyo25; Wirtschafter_Majumder_2023).
@@
Another common type of research study (n=16) examined CCS contributor behavior. While one of these studies used interviews as a data source (yoon25), all of the others relied on the open data from X’s Community Notes. The studies asked descriptive questions about how contributors to the system behaved, such as what types of posts (pilarskiCommunityNotesVs2023; cikm22) and accounts (allen22) contributors tend to leave notes on, what type of content contributors tend to include in notes (jones22; kuuse25), and what types of notes contributors rate as most helpful (solovevReferencesUnbiasedSources2025a; Phillips2025). Three key findings from these studies are that 1) the crowd exerts political bias in the posts that they select to write notes for and the ratings that they assign to notes (allen22); 2) the crowd behaves differently than professional fact-checkers (pilarskiCommunityNotesVs2023), yet often links to them as sources (borenstein2025); and 3) certain types of notes are perceived as more helpful than others, such as those that support claims with links to unbiased (i.e., assessed as politically neutral by a third-party organization [^7] https://mediabiasfactcheck.com/) external sources (solovevReferencesUnbiasedSources2025a).
@@
A third group were design studies that either proposed or built a variation of a CCS (n=10). These papers focused on ways to improve the quantity (liScalingHumanJudgment2025a) and quality (mohammadiAIFeedbackEnhances2025a) of notes, as well as techniques to improve the performance of the note curation algorithm (hawkeye). The bridging-based algorithm for selecting “helpful” notes was introduced in one of these papers (wojcikBirdwatchCrowdWisdom2022b), as a solution to problems of bias (allen22) in a majority-rule note selection system. Recent papers have proposed ways of incorporating AI into the note-writing process and found that LLMs have the potential to generate helpful notes (supernotes; mohammadiAIFeedbackEnhances2025a).
@@
A final group of studies sought to understand user’s perceptions of the notes produced by CCS (n=6). These studies used experiments (liu23), interviews (shusas24), and analysis of X’s data (Hameleers18052024) to answer descriptive questions about how individuals perceived the credibility and helpfulness of context provided by CCS. They found evidence that users perceive notes from CCS as just as, or even more, credible than those from professional fact checkers (liu23; shusas24). While these studies begin to explore the impacts of X’s design decisions, much is still unexplored. As we show next, the potential design space is much wider than what is covered in existing research.
@@ 7. A Design Space for CCS
Our analysis of the various CCS implementations revealed areas of common characteristics shared between the systems as well as points for variation. Analyzing these features allows us to construct a design space of possibilities. Within this design space, we identify six key aspects of CCS:
  1. (1)
    Participation: Who uses the system and what are different users are allowed to do?
  2. (2)
    Inputs: What shape do contributions to the system take?
  3. (3)
    Curation: How are suggested notes selected to be displayed?
  4. (4)
    Presentation: How are selected notes displayed?
  5. (5)
    Platform Treatment: How does the platform treat posts that have displayed notes?
  6. (6)
    Transparency: How can the public observe and evaluate the functioning of the system?
Within each aspect, we identify several dimensions (i.e., points of potential variation), which we frame as design questions, and options (i.e., potential answers for each dimension) based on the features that we observed in our analysis. In the tables in the sections that follow, we indicate when we found evidence that a platform implemented a certain option. However, platforms did not always document all of their design choices, so our tables may be incomplete. In constructing the design space, we additionally include several speculative options that we did not see in any current implementations. The sections that follow detail the dimensions and options for each aspect.
7.1. Participation
@@
Table 2. Participation dimensions and options, as well as platforms that have made this design choice.:

### v3

@@ 5.1. Systematic Literature Review
Figure 3. PRISMA flow diagram (Page et al., 2021) for our systematic literature review, showing both the database search and the backwards reference search. We analyzed a final set of papers (n=56) that combined the results of these two searches.
A PRISMA flow diagram showing the systematic literature review process. It shows two flows side by side: the left is the forwards database search and the right is the backwards reference search. For each flow, the image shows the filtering stages, through identification, screening, and eligibility phases, ultimately resulting in 40 studies from the forward search and 16 from the backwards search included in the qualitative synthesis. Each stage shows the number of records and reasons for exclusions.
In order to synthesize the research community’s understanding of CCS, we performed a PRISMA-style systematic literature review (Page et al., 2021). We sourced relevant papers using both a database keyword search and a backwards reference search (Jalali and Wohlin, 2012). Figure 3 shows a flow diagram summarizing our search and screening process. Our goal was to identify original research that directly studies CCS implementations. Since this category of system has not been consistently named prior to our work, we focused our review on studies of the first large-scale implementation of a CCS by a major social media platform: Twitter’s Community Notes/Birdwatch. We thus used “Birdwatch” and “Community Notes” as keywords and restricted our search to articles that have been written since 2021, when Twitter first launched their implementation. Note that since this is an emerging technology and the set of articles is limited, we decided to include sources of knowledge that are valued in the HCI community but are sometimes excluded from literature reviews, such as pre-prints and non-full-length conference submissions. We thus settled on the following inclusion and exclusion criteria:
@@
Exclusion criteria: the article is not in English, does not present original research, is only an abstract/a summary of other research, or is not about the functioning of a Twitter-style CCS.
@@
We conducted our database search in November 2025. We searched seven databases that are well-known in the fields of HCI and social media research. We began with six that were used in another recent HCI literature review of social media research (Zhang et al., 2023): ACM Digital Library, IEEE Xplore, SAGE Journals, ScienceDirect (Elsevier), Taylor & Francis Online (Tandfonline), and Emerald Insight. We additionally searched arXiv in order to include preprints, which are common in these fields. We searched the full-text, title, abstract, and metadata of articles using the following case-insensitive query string, and limited our results to articles published in 2021 or later:
@@
This search produced n=590 results.
@@
We first screened the search results by title. One researcher read the title for each result and removed articles that were either duplicates or were obviously off-topic, such as those that were not about social media systems. This step resulted in n=254 remaining articles.
@@
We next screened the remaining articles by abstract. One researcher read the article abstracts and removed redundant articles that only summarized research contained elsewhere in our sample, such as books and book chapters, editorials, abstracts, keynotes, and complete conference proceedings. We also removed articles that were topically irrelevant, such as those that did not focus on a Twitter-style CCS, or that used data from such a system, but only to answer research questions unrelated to the system. This step resulted in n=47 remaining articles.
@@
Finally, two researchers read the full text of the remaining articles and screened them further using the inclusion and exclusion criteria. We met during the process to compare results and discuss any differences or challenging cases until we were in agreement. This step produced a set of n=40 articles.
@@
We then used these 40 articles as the starting set for a backwards reference search and extracted all contained references (n=2,207). We de-duplicated these references by paper title, resulting in n=1,385 unique papers. We then performed the same screening steps used in the forward search: we screened titles, which resulted in n=175 papers, then abstracts, which resulted in n=23 papers, then full texts. This backwards search identified n=16 relevant papers not found in the forward search. We combined the sets from both searches into a final set of n=56 papers.
@@
Two researchers then extracted data from each paper in the final set. We agreed upon data to extract, including data source, research questions, main findings, and design implications. We independently read and extracted information from each article, then compared values, discussed disagreements, and reached consensus.
Next, we independently grouped similar papers into affinity groups (Beyer and Holtzblatt, 1999) based on each study’s primary focus. We then compared groupings and discussed our rationale, shifting several papers between groups based on this discussion. Once we agreed upon the groups, we labeled each based on study focus (see Table 1). We describe the results of this analysis in Section 6.
@@ 5.2. Design Feature Identification
In order to identify the design decisions that shape CCS, two authors analyzed four real-world CCS implementations: Twitter/X’s Community Notes [^4] https://x.com/i/communitynotes/, Meta’s Community Notes [^5] https://transparency.meta.com/features/community-notes (which covers Facebook, Instagram, and Threads), TikTok’s Footnotes [^6] https://support.TikTok.com/en/using-TikTok/exploring-videos/footnotes, and YouTube’s Notes [^7] https://support.google.com/youtube/answer/14925346?hl=en. Our analysis was primarily based on public, company-written documentation and communications about the systems and their features, retrieved in July 2025, which we collectively refer to as articles. While we attempted to become contributors for all platforms, some (TikTok, YouTube) did not grant us access. For those that did, this access complemented our understanding and supported our analysis.
Study Focus Primary Method Papers
System Impact (n=24) Data Analysis (X Community Notes Data) (Bouchaud and Ramaciotti, 2025; Renault et al., 2024; Chuai et al., 2025a; Truong et al., 2025; Slaughter et al., 2025; Chujyo et al., 2025; Chuai et al., 2024a, b; Kim et al., 2025; Drolsbach and Pröllochs, 2023; Chuai et al., 2025b; Matamoros-Fernández and Jude, 2025; Razuvayevskaya et al., 2025; Kangur et al., 2024; Braga et al., 2025; Gao et al., 2024; Allen et al., 2024; Bobek and Pröllochs, 2025; Borwankar et al., 2022; Wirtschafter and Majumder, 2023; Wang and Lucas, 2024; Bak-Coleman, 2023; de Keulenaar, 2025; Benjamin, 2022)
Contributor Behavior (n=16) Data Analysis (X Community Notes Data) (Allen et al., 2022; Borenstein et al., 2025; Pilarski et al., 2023; pröllochs2021communitybasedfactcheckingtwittersbirdwatch; Saeed et al., 2022; Kuuse et al., 2025; Jones et al., 2022; Solovev and Pröllochs, 2025; Arjmandi-Lari et al., 2025; Toyoda et al., 2025; Phillips et al., 2025; Renault et al., 2025; Fraxanet et al., 2024; Martel et al., 2025; Mohammadi et al., 2025; Yoon et al., 2025)
System Design (n=10) User studies, design specs, algorithm analysis (Mohammadi and Yasseri, 2025; Wu et al., 2025; Wojcik et al., 2022; Mujumdar and Kumar, 2022; Franzmeyer et al., 2024; Li et al., 2025; De et al., 2025; Zhang et al., 2025; Singh et al., 2025; Xing et al., 2025)
Users’ Perceptions (n=6) User studies (Sharevski et al., 2025; Liu et al., 2023; Shusas and Forte, 2024; Drolsbach et al., 2024; Kankham and Hou, 2025; Hameleers, 2024)
Table 1. The n=56 articles found in our literature review, grouped by study focus.
Summary of existing research categorized by study focus, showing four main categories: System Impact (10 studies using data analysis of X Community Notes data), Contributor Behavior (9 studies using data analysis), User’s Perceptions (6 studies using user studies), and System Design (5 studies using various methods including user studies and algorithm analysis). Each row includes sample paper titles.
For each platform, we independently read the related articles and extracted design details into a spreadsheet to produce a set of design features. We then convened and performed a round of affinity diagramming (Beyer and Holtzblatt, 1999), in which we wrote the design features on sticky notes and placed them on a whiteboard. We placed similar design features together, and through this inductive method identified feature groups. We then used thematic analysis (Braun and Clarke, 2006) to identify the themes of the groupings and came up with provisional group names. We repeated this process with each platform, reconvening to do another round of affinity diagramming after each round of feature extraction. Finally, we synthesized these design features and feature groups into the aspects, dimensions, and options of the design space, which we present below in Section 7.
@@ 6. What Does the Research Community Know About Twitter-Style CCS?
Our systematic literature review of Twitter-style CCS research found n=56 papers and identified four high-level categories of study: studies of the system’s impact, of contributor behavior, of alternative system designs, and of users’ perceptions of crowdsourced context. These categories naturally overlap and are not mutually exclusive, but we highlight them to emphasize the different types of knowledge emerging from this body of research. Table 1 groups the papers according to their most significant contribution, though several papers make multiple contributions and thus span categories.
@@
The most common study type explored CCS impact on the implementing platform (n=24 papers). These studies all analyzed data from X’s Community Notes, asking both causal questions—such as the effect of displaying notes on post diffusion (Drolsbach and Pröllochs, 2023; Chuai et al., 2024b)—and descriptive questions about how X’s CCS performed in specific settings, such as during natural disasters (Chujyo et al., 2025; Bouchaud and Ramaciotti, 2025). Studies found evidence that CCS can reduce both the spread (Drolsbach and Pröllochs, 2023) and believability (Wojcik et al., 2022) of misleading posts annotated with notes. Several articles examined the bridging algorithm used in X’s Community Notes, which is designed to identify high-quality crowdsourced notes likely to be perceived as helpful by users across an ideological spectrum (Bouchaud and Ramaciotti, 2025). Studies showed that while bridging can be effective (Bouchaud and Ramaciotti, 2025), this approach takes significant time to identify “helpful” notes, often causing substantial delays before notes are displayed (Renault et al., 2024; Chuai et al., 2024a, b). Further, studies have shown that the vast majority of suggested notes ($\sim 90\%$) are never published due to insufficient ratings or lack of consensus  (Chujyo et al., 2025; Wirtschafter and Majumder, 2023).
@@
Another common study type (n=16) examined CCS contributor behavior. While one study used interviews (Yoon et al., 2025), all others analyzed open data from X’s Community Notes. Studies asked descriptive questions about contributor actions, such as what types of posts (Pilarski et al., 2023; Saeed et al., 2022) and accounts (Allen et al., 2022) they add context to, what content they include in notes (Jones et al., 2022; Kuuse et al., 2025), and what types of notes they rate as most helpful (Solovev and Pröllochs, 2025; Phillips et al., 2025). Three key findings emerged: 1) contributors show political bias in post selection and note ratings (Allen et al., 2022); 2) contributors behave differently than professional fact-checkers—for example, by focusing on larger accounts (Pilarski et al., 2023)— yet often cite them as sources (Borenstein et al., 2025); and 3) contributors perceive certain note types as more helpful than others. For example, notes that support claims with links to unbiased external sources (i.e., sources assessed as politically neutral by a third-party organization [^8] https://mediabiasfactcheck.com/)  (Solovev and Pröllochs, 2025) are rated more helpful than those that do not.
@@
A third study type was design studies that either proposed or built CCS variations (n=10). These papers focused on ways to improve the quantity (Li et al., 2025) and quality (Mohammadi and Yasseri, 2025) of notes, as well as techniques to improve the performance of the note curation algorithm (Mujumdar and Kumar, 2022). The bridging algorithm for selecting “helpful” notes was introduced in one of these papers (Wojcik et al., 2022), as a solution to bias (Allen et al., 2022) in a majority-rule note selection system. Recent papers have proposed ways of incorporating AI into the note-writing process and found that LLMs have the potential to generate helpful notes (De et al., 2025; Mohammadi and Yasseri, 2025).
@@
A final study type sought to understand users’ (not CCS contributors as above) perceptions of the notes published by CCS (n=6). These studies used experiments (Liu et al., 2023), interviews (Shusas and Forte, 2024), and analysis of X’s data (Hameleers, 2024) to answer descriptive questions about how individuals perceived the credibility and helpfulness of notes published by CCS. These studies found evidence that users perceive CCS notes as just as, or even more, credible than context from professional fact checkers (Liu et al., 2023; Shusas and Forte, 2024).
While these studies begin to explore X’s CCS implementation and its impact, much remains unexplored. As we show next, the CCS design space is potentially much wider than demonstrated by existing research.
@@ 7. A Design Space for CCS
Table 2. Participation dimensions and options, as well as platforms that documented these design choices.
Participation dimensions and options for CCS design, showing 9 dimensions including who sees published notes, who can be contributors, eligibility requirements, contributor limits, incentives, and solicitation methods. Each dimension lists multiple implementation options with corresponding platforms that have adopted each approach
Our analysis of the various CCS implementations revealed common characteristics shared between the systems as well as places for variation. Analyzing CCS features allows us to construct a design space of possibilities for these systems. Within this design space, we identify six key aspects of CCS:
  1. (1)
    Participation: Who can use the system and what are different users allowed to do?
  2. (2)
    Inputs: What form do contributions to the system take?
  3. (3)
    Curation: How are suggested notes selected for display?
  4. (4)
    Presentation: How are selected notes displayed?
  5. (5)
    Platform Treatment: How does the platform treat posts with displayed notes?
  6. (6)
    Transparency: How may the public observe and evaluate the system’s functioning?
Within each aspect, we identify several dimensions (i.e., points of potential variation), which we frame as design questions, and options (i.e., potential answers for each dimension) based on the features observed in our analysis. In the tables in the sections that follow, we indicate when we found evidence that a platform implemented a certain option. However, platforms did not document all of their design choices, so our representations of each platforms’ features should be seen as illustrative, not definitive. In constructing the design space, we additionally include several speculative options not seen in current implementations, but which demonstrate important potential variations. The sections that follow detail the dimensions and options for each aspect.
7.1. Participation